The "Home" country has 1200 units of labor at its disposal. It can produce two goods: apples
and bananas. To produce 1 kg of apples, 3 units of labor are required, for 1 kg of bananas, 2
units of labor are required.
a) Graphically represent the production options curve of Home.
b) What is the opportunity cost for the Home of apples in terms of bananas?
c) In the absence of trade, what will be the price of apples in terms of bananas?

Now imagine that there is another country, "Outside," which has 800 units of labor at its
disposal. It also produces apples and bananas. To produce 1 kg of apples, 5 units of labor are
required, for 1 kg of bananas, 1 unit of labor is required.
d) Draw the "Outside" production options curve on a graph.
e) Suppose that in autarkie (= self-sufficient economy), each country uses half of its labor
supply in the production of each product. What is the production? Present this on the graph in
(d).
f) What would be a trade situation (specialization + terms of trade) that would be beneficial
for both countries?
g) Between which two values should the terms of trade lie so that trade is beneficial for both
countries?
Answer #1

a)

In case of country Home

Maximum production of apples=1200/3=400 unit of apples

Maximum production of banana=1200/2=600 unit of banana

Please refer graph given at the bottom of solution.

b)

In case of country Home

Opportunity cost of one unit of apples=600/400=1.50 unit of banana

c)

In case of country Home

Price of apple=opportunity cost of apple=1.50 units of banana/unit of apple

In case of outside

d)

In case of country Outside

Maximum production of apples=800/5=160 unit of apples

Maximum production of banana=800/1=800 unit of banana

Please refer graph given at the bottom of solution.

e)

In case of home

Production of apples=600/3=200 unit of apples

Production of banana=600/2=300 unit of banana

In case of country Outside

Production of apples=400/5=80 unit of apples

Production of banana=400/1=400 unit of banana

f)

Opportunity cost of one unit of apples in country outside=800/160=5 units of banana

We can see that opportunity cost of apples is lower in case of Home. So,

Home should specialize in production of apples and Outside should specialize in production of banana.

Home can produce apples and keep half of the production for internal consumption, and export the rest to outside.Outside can produce banana and keep half off the production for internal consumption and export the rest to Home.

Terms of trade should for apple (say) should be in between the autarky prices (opportunity cost in each country)

g)

Trade would be beneficial if terms of trade is between 1.50 units of banana and 5 units of banana for 1 unit of apples